Hi, I am struggling with the whole alegbra concept so have two questions:
1. Eight washington stamps and two lincoln stamps cost $2.80; three washingston stamps and four lincoln stamps cost $2.35; Find the cost of each stamp and show the solution equations?

2. One number is 18 more than another, and the sum of the smaller number is twice the larger number is 45. What are the two numbers.
If someone could help I would be very happy. Thank you.

Answer by stanbon(48574)

1. Eight washington stamps and two lincoln stamps cost $2.80; three washingston stamps and four lincoln stamps cost $2.35; Find the cost of each stamp and show the solution equations?
Let W be the cost of the Washington stamps and L be the cost of the Lincoln.
EQUATIONS:
1st: 8W + 2L=280 cents
2nd: 3W + 4L=235 cents
Multiply the 1st equation by 2 to get:
3rd: 16W + 4L = 560
Subtract the 2nd equation from the 3rd to get:
4th: 13W=325
W= 25 cents (this is the cost of each Washington stamp)
Substitute this value in 1st or 2nd to find L, as follows:
3(25)+4L=235
75+4L=235
4L=160
L=40 cents (this is the cost of each Lincoln stamp)
-------------------------------------------------------
2. One number is 18 more than another, and the sum of the smaller number and twice the larger number is 45. What are the two numbers.
Let the smaller number be "x".
Then the larger number is "x+18"
EQUATION:
smaller + 2(larger) = 45
x + 2(x+18) = 45
x + 2x + 36 = 45
3x = 9
x = 3 (this is the smaller number)
x+18 = 21 (this is the larger number)
------------------------------------------------
Cheers,
Stan H.